# Demystifying Quantum Power Flow: Unveiling the Limits of Practical Quantum Advantage

### v2

@@ IV.4 End-to-End Complexity
Assuming access to a QRAM pre-loaded with $|\mathbf{b}\rangle$ the most optimistic complexity estimate becomes
@@
Now, as $\kappa\gg 1$ from Figure 3 (center) we get the leading order, asymptotic complexity as (considering the best case for PQA in QPF)
@@
It is important to understand that (4) does not represent the actual runtime of solving DCPF using the HHL algorithm. Rather it gives the asymptotic scaling of this runtime with the parameters of the problem. This scaling is independent of the specific hardware and algorithmic implementations used, while the exact runtime will depend on those factors.
@@ V Discussion and Conclusion
This section presents a comparative discussion between CG and end-to-end quantum complexity from PQA view point. In Table 1, end-to-end complexities of quantum methods are compared with CG. We find that an additional $\log(N)$ factor is required in the CG complexity to make the error guarantees of CG and HHL comparable (see supplementary information for more details). We see that due to state preparation, readout (tomography) complexity, and condition number scaling, quantum algorithms have higher complexity than CG. This negates any chance of PQA in DCPF in this setting unless fundamental algorithmic barriers are overcome. Further, this end-to-end complexity comparison can directly be extended to the FDLF formulation presented in [23]. Even with the presence of QRAM, quantum computing’s run-time will be asymptotically worse with system size for FDLF solving compared to CG, even if FDLF takes $\mathcal{O}(\log(N))$ iterations to converge classically. [^9] With linear system solving iterations growing as $\mathcal{O}(\log(N))$, FDLF can be solved classically with complexity $\mathcal{O}(N\log(N)\sqrt{\kappa}s\log(1/\varepsilon))$ which is better in terms of error, sparsity and condition number from HHL complexity of the same with QRAM. Therefore, it can be concluded that practical quantum advantage does not exist for solving power flow problems (both DCLF and FDLF), primarily due to the requirement of reading out the complete solution vector as well as due to the complexity of encoding the injection vector into a quantum state. Figure 5 (left) illustrates the runtime complexity comparisons using actual power network data (as shown in Figure 3) and highlights that CG runtime complexity scales $\sim N^{2.5}$ which is significantly lower compared to HHL’s runtime complexity growth $\sim N^{5.5}$ and VTAA-HHL’s $\sim N^{4.5}$.
@@
One of the major computational bottlenecks in the quantum approach is the readout complexity which scales linearly with $O(N).$ This can be reduced if we only need to read $D$ components of the solution vector instead of all the $N$ components. We call this the readout level of the problem and assume that $D\ll N$. This can reduce the readout complexity to $O(D).$ From our earlier complexity analysis, we can formulate a relation between the various parameters in the problem ($D,s,\kappa,\ldots$) such that PQA becomes possible. As complexities do not represent exact runtimes, for there to be PQA the complexity of CG must be greater (e.g. a factor of $N^{\alpha}$ where $\alpha>0$) than that of a quantum method. For the optimal HHL complexity from Table 1 to be substantially less than the complexity of CG, we require that $D\kappa^{3/2}\ll Ns^{-1}.$ Similarly, in the optimistic scenario with the VTAA-HHL algorithm, we get $D\sqrt{\kappa}\log^{3}(\kappa)\ll Ns^{-1}.$ These relations act as a rule of thumb to check for the existence of PQA in these problems.
Complexity scaling arguments do not preclude the possibility that a single instance of DCPF may be solved faster on a quantum machine in comparison to CG. However, this perceived advantage will eventually go away as the problem gets more computationally intense due to an increase in system size or changes in other parameters. Moreover, the fact that a single instance can be solved faster does not imply a fundamental quantum speed-up and only points to differences in hardware and implementation.
@@
Figure 4: Composition of End-to-End complexity of solving QPF using HHL. Here, $T_{b}$ represent state preparation runtime cost in terms of circuit depth, $T_{s}$ is cost of a single quantum simulation to solve linear system, and $T_{r}$ represent the number of copies of $|\tilde{{\mathbf{x}}}\rangle$ needed to obtain $\widehat{\mathbf{x}}$ with a given error. With QRAM $T_{b}=\log(N)$ the end-to-end complexity is $\mathcal{O}\big{(}N\log(N)\kappa^{2}s^{2}/\varepsilon^{2}\big{)}$ and if QRAM is not available, than arbitrary state preparation complexity $T_{b}=N$ need to be considered which leads to end-to-end complexity as $\mathcal{O}\big{(}N^{2}\kappa/\varepsilon+N\log(N)\kappa^{2}s^{2}/\varepsilon\big{)}$ for complete $N$ state readout.
@@
Figure 5: Runtime complexity comparison and potential PQA parameter range. Left: Complexity comparison between CG, HHL-Optimistic and VTAA-HHL-Optimistic for power networks from PGLib[35]. The fit beyond the 500-Bus system indicate that approximate end-to-end runtime complexity of CG is $\mathcal{O}(N^{2.5})$, which is significantly lower than HHL’s $\mathcal{O}(N^{5.5})$ and even VATT-HHL’s end-to-end runtime complexity $\mathcal{O}(N^{4.5})$. Therefore, it is evident that there is no practical quantum advantage in solving DCPF for these systems. Right: Upper bound of condition number ($\kappa$) values for achieving PQA using HHL complexity (4), at different readout levels for system size $N$. Dotted black lines indicate different readout levels and $D=10$ means that only 10 classical values need to be extracted from the solution state $|\mathbf{x}\rangle$. Below these curves, the runtime complexity of CG is worse than that of HHL and there is a possibility of PQA. Importantly, for the PGLib systems we cannot achieve any quantum speedup even for a single readout without reducing the condition number. It is clear from Table 1 and related discussion, that PQA can only be achieved if the condition number is reduced using preconditioning and readout requirements are limited to a fraction of the state vector length. Here complexity dependence on error is ignored considering that errors do not scale with system size. Also, we consider machine overheads to be insignificant.
@@
Figure 5 (right) illustrates the interplay between condition number $\kappa$, system size $N$, and readout level $D$ for attaining PQA, at constant sparsity. Notably, higher readout levels demand lower condition numbers for PQA. In a power system, it is a well-known fact that the number of branches reaching or violating desired limits is only a fraction of the total number of branches. Thus, a problem to identify the power flow on candidate violating branches can be cast as a linear system with very low readout level requirements. However, for PQA, this system would still require preconditioning to reduce the condition number. Therefore, exploring techniques to mitigate condition number growth is essential for achieving PQA. Most importantly, the conditions in Figure 5 highlight the possibility of having PQA with at most sub-quadratic speedup. For near-exponential speedup, both the condition number and readout requirements should ideally not scale with the size of the system.
@@
In conclusion, our end-to-end analysis of QPF complexity reveals a nuanced picture of PQA in power flow calculations. While initial claims of exponential speedups proved illusory, this research pinpoints a narrow window where PQA might be achievable. For DCPF and FDLF problems, our findings demonstrate the absence of PQA under typical network parameters and problem specifications. However, for DCPF-type problems with specific condition number values and readout requirements, a sub-quadratic speedup may be attainable. This highlights the need for further research into tailored QPF algorithms and problem settings to unlock the true potential of quantum computing for power grid analysis. This work provides a critical step towards demystifying the hype surrounding QPF and understanding the requirements for future breakthroughs in this field.
Supplementary Information
Demystifying Quantum Power Flow: Unveiling the Limits of Practical Quantum Advantage
@@ B. Brief Explanation of HHL
The algorithm commences by encoding the load vector $|b\rangle$ information onto an $n=\log(N)$ qubit register using a unitary operation $U_{b}$ such that $U_{b}|0\rangle=|\mathbf{b}\rangle$. This action signifies that the initial value register is now in a state containing the information from $|\mathbf{b}\rangle$. Subsequently, the quantum phase estimation subroutine is employed to encode the eigenvalues of matrix $A$ into an $n_{c}$-qubit register. The number of qubits ($n_{c}$) determines the accuracy with which the phase values, i.e., eigenvalues, are stored. The next step involves applying controlled rotation to an ancilla register (one qubit) based on the eigenvalues stored in the $n_{c}$-register. This rotation utilizes the principle that for a Hermitian $A$, the eigenvalues of $A^{-1}$ are the inverses of the eigenvalues of $A$, using the spectral decomposition theorem. In other words, if $A=Q\Lambda Q^{\dagger}$ with $Q$ as the eigenvector matrix and $\Lambda$ as the diagonal matrix with eigenvalue entries, then $A^{-1}=Q\Lambda^{-1}Q^{\dagger}$. Upon rotation, the probability of obtaining $|1\rangle$ in the ancilla register outcome, through measurement, is proportional to the eigenvalues of $A^{-1}$, i.e., the inverses of the eigenvalues of $A$. Then an inverse phase estimation is performed to return the $n_{c}$-qubit register to its base state $|0\rangle$ and finally, the ancilla is measured. Now, if the ancilla measurement result is $|1\rangle$ then the algorithm succeeds and we get the solution as a quantum state. Otherwise, the entire process starting from state preparation to ancilla measurement is repeated until the state $|1\rangle$ is achieved. The final quantum state $|\tilde{\mathbf{x}}\rangle$ is proportional to the solution vector $|\mathbf{x}\rangle$. Readers can refer to [13, 37, 11, 10] for details and explanation of HHL workings.
@@
i) Hot Starting: It is clear that if classical and quantum computing errors are less than a threshold, they have a lesser impact on comparative run-time performance analysis. However, by definition of classical error $\varepsilon_{c}$ it is clear that hot-starting or warm-starting a CG-based DCPF solving algorithm can lead to faster run-times as discussed in the main section. In practical situations, we run multiple DCPFs with load variations within a range. Thus, warm-starting will benefit in reducing the overall time taken by the DCPF running over multiple instances. Although in literature the ideas of warm-starting of quantum algorithms have been presented [43], potential benefits, if any, will require closer scrutiny and dedicated quantum linear system warm-starting methods.
ii) Condition Number limitations: The HHL algorithm (and other variants as well) are shown to be robust when the $A$ matrix is well conditioned i.e. low condition number. However, in a power system, the $A$ matrix condition number grows approximately quadratically with the system size as shown in Figure 3(b). Authors in [23] discuss this limitation on condition number and suggest that pre-conditioning methods can be developed to artificially reduce the condition number of the $A$ matrix. However, there are a few limiting factors which will impact any envisioned benefit from preconditioning:
  • Pre-conditioner selection is not a trivial problem. Also, the complexity of preconditioning will modify the overall complexity of the linear system-solving process. To maintain a practical quantum advantage, preconditioning complexity must be less than that of solving the linear system itself. For example $A^{-1}$ is a pre-conditioner of $A$ as $A^{-1}A=\rm I$. However finding $A^{-1}$ is equivalent of solving the linear system $A\mathbf{x}=\mathbf{b}$. Therefore, it is difficult to establish if PQA will exist for power flow problems, without analyzing specific preconditioning algorithm.
  • Benefits of preconditioning will be available for both classical and quantum algorithms, given that preconditioning complexity is acceptable as discussed in the aforementioned point. Also, note that HHL’s run-time complexity dependence on condition number $\kappa$ is worse than that of CG’s. This implies condition number reduction will impact only the robustness of QPF methods and unlike to play any role in deciding PQA for QPF.

### v3

@@ IV-D End-to-End Complexity
Assuming access to a QRAM pre-loaded with $|\mathbf{b}\rangle$, the most optimistic complexity estimate becomes
@@
As $\kappa\gg 1$ as shown in Figure 3 (center), the leading order, asymptotic complexity (considering the best case for PQA in QPF) is
@@
It is important to understand that (4) does not represent the actual runtime of solving DCPF using the HHL algorithm. Rather, it gives the asymptotic scaling of this runtime with the parameters of the problem. This scaling is independent of the specific hardware and algorithmic implementations used and the exact runtime will depend on those factors.
@@ V Discussion and Conclusion
This section presents a comparative discussion between CG and end-to-end quantum complexity from a PQA viewpoint. In Table I, end-to-end complexities of quantum methods are compared with CG.
Using the same methodology we also report the end-to-end complexities of a more advanced quantum system solver, namely HHL with Variable Time Amplitude Amplification (VTAA-HHL), in Table I. VTAA-HHL was introduced by Ambainis [14]. VTAA-HHL, combines HHL with VTAA, which is a method to boost the success probability of the HHL algorithm with improved complexity. As discussed before, the HHL circuit fails with probability $O(1/\kappa^{2})$. In the vanilla version of HHL this is boosted to a constant success probability by using the amplitude amplification technique. Unlike vanilla amplitude amplification, which assumes a uniform runtime and applies a fixed number of Grover iterations, VTAA adapts the amplification process to selectively boost the success amplitude leading to a linear system solver that runs in time linear in $\kappa$ .
In final row of Table I, we also compute the theoretically optimal end-to-end complexity of a quantum linear system solver running on a fault tolerant quantum computer. This bound is derived from the known complexity lower bound for this problem obtained in the original HHL paper. This calculation necessarily lower bounds the solving complexity of all modern quantum solvers [47]. We see that despite assuming this highly optimistic scenario with a theoretically optimal quantum solver and $O(\log(N))$ data loading overhead, the end-to-end complexity of the quantum solver is still worse than that of CG.
@@
We find that an additional $\log(N)$ factor is required for CG complexity to make the error guarantees of CG and HHL comparable (see appendix for more details). Due to state preparation, readout (tomography) complexity, and condition number scaling, quantum algorithms have higher complexity than CG. This negates any chance of PQA in DCPF in this setting unless fundamental algorithmic barriers are overcome. Further, this end-to-end complexity comparison can directly be extended to the FDLF formulation presented in [24]. Even with the presence of QRAM, quantum computing runtime will be asymptotically worse in system size for FDLF compared to CG, even if FDLF takes $\mathcal{O}(\log(N))$ iterations to converge classically. [^9] With linear system solving iterations growing as $\mathcal{O}(\log(N))$, FDLF can be solved classically with complexity $\mathcal{O}(N\log(N)\sqrt{\kappa}s\log(1/\varepsilon))$, which is better in terms of error, sparsity, and condition number from HHL complexity of the same with QRAM. Therefore, practical quantum advantage does not exist for solving power flow problems (both DCLF and FDLF), primarily because of the requirement of reading out the complete solution vector and because of the complexity of encoding the injection vector into a quantum state. Figure 5 (left) illustrates the runtime complexity comparisons using actual power network data (as shown in Figure 3) and highlights that CG runtime complexity scales $\sim N^{2.5}$, which is significantly lower compared to HHL’s runtime complexity growth $\sim N^{5.5}$ and VTAA-HHL’s $\sim N^{4.5}$.
@@
One major computational bottleneck in the quantum approach is readout complexity, which scales linearly as $O(N).$ This can be reduced for some specific power flow application if only $D$ components of the solution vector need to be read instead of all $N$ components. We call this the readout level of the problem and assume that $D\ll N$. This can reduce the readout complexity to $O(D).$ From our earlier complexity analysis, we formulate a relation between the various parameters in the problem ($D,s,\kappa,\ldots$) such that PQA becomes possible. As complexities do not represent exact runtimes, for there to be PQA, the complexity of CG must be greater (e.g., a factor of $N^{\alpha}$ where $\alpha>0$) than that of a quantum method. For the optimal HHL complexity from Table I to be substantially less than the complexity of CG, $D\kappa^{3/2}\ll Ns^{-1}$ is required. Similarly, in the optimistic scenario with the VTAA-HHL algorithm, we get $D\sqrt{\kappa}\log^{3}(\kappa)\ll Ns^{-1}.$ These relations act as a general rule to check for the existence of PQA in these problems.
Even in a regime where this rule of thumb does not hold, quantum computers may be able to solve specific DCPF instances faster than CG for small systems. But this doesn’t imply a fundamental quantum speedup. Moreover near-term fault tolerant quantum computers are expected to have significant error-correction overheads [48, 49]. Thus, for PQA these overheads should be offset by a significantly favorable complexity scaling for the quantum algorithm. This can further narrow the window of parameters for which PQA can exist.
@@
Figure 4: Composition of end-to-end complexity of solving QPF using HHL. Here, $T_{b}$ represents state preparation runtime cost (in terms of circuit depth), $T_{s}$ is the cost of a single quantum simulation to solve linear system, and $T_{r}$ represents the number of copies of $|\tilde{{\mathbf{x}}}\rangle$ needed to obtain $\widehat{\mathbf{x}}$ with a given error. With QRAM $T_{b}=\log(N)$, the end-to-end complexity is $\mathcal{O}\big(N\log(N)\kappa^{2}s^{2}/\varepsilon^{2}\big)$, and if QRAM is not available, than arbitrary state preparation complexity $T_{b}=N$ needs to be considered, which leads to end-to-end complexity of $\mathcal{O}\big(N^{2}\kappa/\varepsilon+N\log(N)\kappa^{2}s^{2}/\varepsilon\big)$ for complete NN state readout. Readers can refer to [13, 37] for HHL’s quantum circuit and Figure 1 of [24] for quantum circuits for solving QPF.
@@
Figure 5: Runtime complexity comparison and potential PQA parameter range. Left: Complexity comparison between CG, HHL-Optimistic, and VTAA-HHL-Optimistic for power networks from PGLib[34]. The fit beyond the 500-Bus system indicates that the approximate end-to-end runtime complexity of CG is $\mathcal{O}(N^{2.5})$, which is significantly lower than HHL’s $\mathcal{O}(N^{5.5})$ and even VTAA-HHL’s $\mathcal{O}(N^{4.5})$. Therefore, there is no practical quantum advantage in solving DCPF for these systems. Right: Upper bound of condition number ($\kappa$) values for achieving PQA using HHL complexity (4) at different readout levels for system size $N$. Dotted black lines indicate different readout levels and $D=10$ means that only 10 classical values need to be extracted from the solution state $|\mathbf{x}\rangle$. Below these curves, the runtime complexity of CG is worse than that of HHL and there is a possibility of PQA. Importantly, for the PGLib systems, we cannot achieve any quantum speedup even for a single readout without reducing the condition number. It is clear from Table I and related discussion that PQA can only be achieved if the condition number is reduced using preconditioning and readout requirements are limited to a fraction of the state vector length. Here, complexity dependence on error is ignored as errors do not scale with system size. Also, machine overheads are considered insignificant.
@@
Figure 5 (right) illustrates the interplay between condition number $\kappa$, system size $N$, and readout level $D$ for attaining PQA at constant sparsity. Notably, higher readout levels demand lower condition numbers for PQA. In a power system, it is well-known that the number of branches reaching or violating desired limits is only a fraction of the total number of branches. Thus, a problem to identify the power flow on candidate violating branches can be cast as a linear system with very low readout level requirements. However, for PQA, this system would still require preconditioning to reduce the condition number. Therefore, exploring techniques to mitigate condition number growth is essential for achieving PQA. Most importantly, the conditions in Figure 5 highlight the possibility of having PQA with at most sub-quadratic speedup. For near-exponential speedup, both the condition number and readout requirements should ideally not scale with the size of the system.
@@
In conclusion, our end-to-end analysis of QPF complexity reveals a nuanced picture of PQA in power flow calculations. Although initial claims of exponential speedups proved illusory, this research pinpoints a narrow window where PQA might be achievable. For DCPF and FDLF problems, our findings demonstrate the absence of PQA under typical network parameters and problem specifications. However, for DCPF-type problems with specific condition number values and readout requirements, a sub-quadratic speedup may be attainable. This highlights the need for further research into tailored QPF algorithms and problem settings to unlock the true potential of quantum computing for power grid analysis. This work provides a critical step towards demystifying the hype surrounding QPF and understanding the requirements for future breakthroughs in this field.
A. Brief Explanation of the CG Algorithm
CG iteratively finds the minimum of $f(\mathbf{x})=\frac{1}{2}\mathbf{x}^{T}A\mathbf{x}-\mathbf{b}^{T}\mathbf{x}+c$, which is a solution of $A\mathbf{x}=\mathbf{b}$ if $A$ is positive definite. The main idea of CG is to generate vectors at each iteration that are conjugate (or A-orthogonal) to all previous iteration vectors and only require the $i$-th conjugate vector for generating the $i+1$-th conjugate vector[9]. Using the conjugate property of these vectors (called direction vectors), CG is shown to require at most $N$ iterations to converge to the solution of $A\mathbf{x}=\mathbf{b}$, from any starting point $\mathbf{x}^{o}\in\mathbb{R}^{N}$. However, the floating point round-off errors and cancellation errors prevent using CG to find an exact solution. Also, for DCPF problems, we are interested in finding a solution while accepting $\texttt{error}\leq\texttt{Measuring Device Precision}$. In performing these conjugate vector calculations and updated states in CG, matrix-vector multiplication becomes the dominant operation [9]. For a sparse matrix with at max $s$ non-zero elements in a row, one matrix-vector multiplications will have complexity $\mathcal{O}(Ns)$.
The second part of CG complexity depends on the number of these matrix-vector multiplications to find the solution $\mathbf{x}^{\star}$ with an acceptable error $\varepsilon_{c}$. This means that the number of iterations required can be calculated based on error convergence after $i$-iterations, i.e., the relationship between initial- and post-$i$-th iteration error. Let initial error vector $\mathbf{e}^{o}=\mathbf{x}^{\star}-\mathbf{x}^{o}$, and after $i$ iterations, $\mathbf{e}^{i}=\mathbf{x}^{\star}-\mathbf{x}^{i}$ then error convergence will be $\|\mathbf{e}^{i}\|_{A}\leq\varepsilon_{c}\|\mathbf{e}^{o}\|_{A}$. Here, $\|\mathbf{e}\|^{2}_{A}=\mathbf{e}^{T}A\mathbf{e}$ is sometimes referred to as energy-norm [9] and $\varepsilon_{c}$ is the error constant that determines how close CG convergence is to the actual solution with respect to the initial error ($\varepsilon_{c}\geq\|\mathbf{e}^{i}\|_{A}\big/\|\mathbf{e}^{o}\|_{A}$ ). Further, using convergence analysis from [9], the maximum number of iterations needed to achieve $\varepsilon_{c}$ is given as
$i\leq\bigg\lceil\frac{1}{2}\sqrt{\kappa}\log(2/{\varepsilon_{c}})\bigg\rceil$
The square root dependence of CG runtime complexity on condition number $\kappa$ is beneficial because a power system network matrix $A$ tends to have a high condition number and approximately quadratic growth of $\kappa$ with respect to system size $N$ for PGLib networks. The feature of power system matrices having high condition numbers affects the convergence of CG as a high condition number and implies a higher eigenvalue spread, meaning a large number of iterations are required to converge [9]. Preconditioning of the linear system $A\mathbf{x}=\mathbf{b}$ can solve this problem. However, finding a computationally cheap and stable preconditioner is a non-trivial problem [9].
For comparing the complexity of CG with HHL-QPF, it’s crucial to establish the relationship between $\varepsilon_{c}$ (error defined using the energy norm) and $\varepsilon$ (error defined using the Euclidean norm). If $A$ is positive semidefinite (PSD) and $\lambda_{1}<\dots<\lambda_{N}$ represents the ordered set of eigenvalues of $A$, then $\lambda_{1}\varepsilon\leq\varepsilon_{c}$. Thus, running CG until the energy norm of the error is less than $\varepsilon/\lambda_{1}$ guarantees that the $L_{2}$ error in the solution is less than $\varepsilon$. This unifies the error guarantees of CG and HHL and brings their complexities onto the same footing. Additionally, as shown in Figure 6, for PGLib systems, the smallest eigenvalue grows approximately as $N^{x}$ with $N$ being the number of buses. Consequently, the CG complexity in Table I will include a $\log(N)$ term when $\varepsilon_{c}$ with $\varepsilon$ is substituted in the end-to-end complexity.
@@ B. Brief Explanation of HHL
The algorithm starts by encoding load vector $|b\rangle$ information onto an $n=\log(N)$ qubit register using a unitary operation $U_{b}$, such that $U_{b}|0\rangle=|\mathbf{b}\rangle$. This action signifies that the initial value register is now in a state containing the information from $|\mathbf{b}\rangle$. Subsequently, the quantum phase estimation subroutine is employed to encode the eigenvalues of matrix $A$ into an $n_{c}$-qubit register. The number of qubits ($n_{c}$) determines the accuracy with which the phase values, i.e., eigenvalues, are stored. The next step involves applying controlled rotation to an ancilla register (one qubit) based on the eigenvalues stored in the $n_{c}$-register. This rotation uses the principle that for a Hermitian $A$, the eigenvalues of $A^{-1}$ are the inverses of the eigenvalues of $A$, using the spectral decomposition theorem. In other words, if $A=Q\Lambda Q^{\dagger}$ with $Q$ as the eigenvector matrix and $\Lambda$ as the diagonal matrix with eigenvalue entries, then $A^{-1}=Q\Lambda^{-1}Q^{\dagger}$. Upon rotation, the probability of obtaining $|1\rangle$ in the ancilla register outcome, through measurement, is proportional to the eigenvalues of $A^{-1}$, i.e., the inverses of the eigenvalues of $A$. Then an inverse phase estimation is performed to return the $n_{c}$-qubit register to its base state $|0\rangle$, and finally, the ancilla is measured. If the ancilla measurement result is $|1\rangle$, then the algorithm succeeds and the solution is provided as a quantum state. Otherwise, the entire process starting from state preparation to ancilla measurement is repeated until the state $|1\rangle$ is achieved. The final quantum state $|\tilde{\mathbf{x}}\rangle$ is proportional to the solution vector $|\mathbf{x}\rangle$. Readers can refer to [13, 37, 11, 10] for details and an explanation of how HHL works.
@@
(i) Hot-starting: It is clear that if classical and quantum computing errors are less than a threshold, they have a lesser impact on comparative runtime performance analysis. However, by the definition of classical error, $\varepsilon_{c}$, it is clear that hot-starting or warm-starting a CG-based DCPF solving algorithm can lead to faster runtimes as discussed in the main section. In practical situations, we run multiple DCPFs with load variations within a range. Thus, warm-starting will benefit in reducing the overall time taken by the DCPF running over multiple instances. Although the idea of warm-starting quantum algorithms has been presented in literature [50], potential benefits, if any, will require closer scrutiny and dedicated quantum linear system warm-starting methods.
(ii) Condition number limitations: The HHL algorithm (and other variants) are shown to be robust when the $A$ matrix is well conditioned, i.e., has a low condition number. However, in a power system, the $A$ matrix condition number grows approximately quadratically with the system size, as shown in Figure 3(center). Authors in [24] discuss this limitation on condition number and suggest that preconditioning methods can be developed to artificially reduce the condition number of the $A$ matrix. However, there are a few limiting factors that will impact any envisioned benefit from preconditioning:
  • Preconditioner selection is not a trivial problem. Also, the complexity of preconditioning will modify the overall complexity of the linear system-solving process. To maintain a practical quantum advantage, preconditioning complexity must be less than that of solving the linear system itself. For example $A^{-1}$ is a preconditioner of $A$ as $A^{-1}A=\rm I$. However finding $A^{-1}$ is equivalent to solving the linear system $A\mathbf{x}=\mathbf{b}$. Therefore, it is difficult to establish if PQA will exist for power flow problems without analyzing specific preconditioning algorithms.
  • Benefits of preconditioning will be available for both classical and quantum algorithms given that preconditioning complexity is acceptable as discussed prior. Also, note that HHL’s runtime complexity dependence on condition number $\kappa$ is worse than that of CG’s. This implies that condition number reduction will impact only the robustness of QPF methods and is unlikely to play any role in deciding PQA for QPF.